# TraCE: Trajectory Counterfactual Explanation Scores

### v1

@@ 4.2.2 Results and Discussion
Analysis of the SSP TraCE scores for different countries found that most countries in the study fit a common pattern (Figure 3). Figure 4 shows the cumulative TraCE score time series (2015-2022) for Norway, which was identified as a representative country. The TraCE score trajectories are consistently positive across all SSPs, in agreement with the expectation that they were developed as realistic scenarios in alignment with the factual historical data. Of note is the visible flattening around the year 2020, which coincides with the onset of the COVID-19 pandemic. This flattening occurs because the SSP projections did not anticipate the pandemic, so the observed data features deviate from these trajectories, resulting in low or negative instantaneous TraCE scores. Overall, Figure 4 indicates that SSP5 (Fossil-fueled Development) consistently ranks the highest from late 2015 onwards, while other SSPs score more closely together. However, starting in mid-2021, the SSP4 (Inequality) and SSP1 (Sustainability) TraCE scores begin to diverge above those of SSP2 (Middle of the Road) and SSP3 (Regional Rivalry).
A heatmap is shown in Figure 3 for the average TraCE scores of 15 countries in the study. This provides an overview of the countries’ alignments with SSP projections for the study period 2015-2022. Comparisons can be made between SSPs for a single country, and across different countries. A common pattern emerges across most countries (as visualised in Figure 4 for Norway), with SSP5 ranking highest, followed by SSP1, closely tracked by SSP4, and finally, SSP2 and SSP3 ranking closely together. Some notable results stand out: several countries, including Germany, Greece, Italy, Mexico, and Portugal, exhibit lower TraCE scores across all SSPs. This indicates that their observed data features are less similar to their corresponding SSP projections, when compared to most other countries in the study. Additionally, some countries deviate from the majority SSP ranking pattern. For example, Greece aligns most closely with SSP4, followed by SSP2 and SSP3, with SSP1 and SSP5 ranking the lowest. Italy aligns most with SSP3, showing strong divergence from the remaining SSPs, which have similar TraCE scores. Poland closely aligns with SSP4, followed by SSP3, with TraCE scores diverging significantly from the other SSPs. Additional plots presenting the findings for Poland, as an example, are available in Appendix A.3.
Figure 3: Country heatmap of average TraCE scores for each SSP, for the period 2015-2022. Higher TraCE score indicates closer alignment with a Shared Socioeconomic Pathway (SSP).
The utility of TraCE scores in this application lies in the capability to reconcile complex and occasionally conflicting variables into a single value. This allows non-experts to quickly assess alignment with the established SSP scenarios, via an explainable method based on direction and distance in the data feature space. Visually assessing such alignment from the raw data itself can be challenging, particularly as the number of included features increases. The TraCE method is therefore useful for communication and understanding between stakeholder groups.
It must be emphasised that this study serves as a proof of concept, and requires input from experts across multiple domains to ensure safe and trustworthy downstream implementation. This includes the selection of data features for monitoring, and their weighting (which has been equally distributed in this demonstration). Different weighting schemes will yield distinct results and should be developed in accordance with the priorities and specific questions of the user. Additionally, the data used and results obtained are contingent on the model source for SSP projections.
@@ 5 Conclusion
TraCE provides a model-agnostic modular framework from which to assess progress over time towards an assigned goal. As demonstrated, the modularity of TraCE enables application-specific adaptation. For example, target points can be defined as most appropriate: model-generated counterfactuals, curated points hand selected by experts or established industry benchmarks. Deployment and interpretation of TraCE should be guided by domain experts and the implementation of TraCE for the presented applications are for illustrative purposes only. By distilling high dimensional dynamic sequential tasks into a single value, TraCE scores enable experts and laypeople alike to quantify and better understand progress in sequential tasks.
@@ A.2 Intensive care unit outcomes
(a) Instantaneous TraCE scores: higher indicates more alignment with a counterfactual.
(b) Classifier probabilities via prediction model.
Figure 6: Patient journey for a unsuccessfully discharged patient (in-hospital mortality). NRFD = Not ready for discharge, RFD = ready for discharge (desirable outcome), mortality (undesirable outcome).
In contrast to the presented ICU stay which concluded in successful discharge (Figure 1(a)), here we present a negative outcome ICU stay which resulted in in-hospital mortality. For the first half of the stay the classifier predicts that the time point is most likely to be not ready for discharge (NRFD) closely followed by mortality (Figure 5(b). The high mortality probability is reflected in the instantaneous TraCE score which aligns more with the undesirable (mortality) counterfactual than the desirable (ready for discharge), and negative trend in total TraCE score (Figure 5(a). This suggests that TraCE may be suitable for not just early detection of patient recovery but also deterioration.

### v2

@@ 4.2.2 Results and Discussion
Analysis of the average TraCE scores for 15 different countries found that most countries in the study fit a common pattern. An overview of the countries’ alignments with SSP projections is shown in the heatmap (Figure 3) for the study period 2015-2022. Comparisons can be made between SSPs for a single country, and across different countries. A common pattern emerges across most countries, with SSP5 (Fossil-fueled Development) ranking highest, followed by SSP1 (Sustainability), closely tracked by SSP4 (Inequality), and finally, SSP2 (Middle of the Road) and SSP3 (Regional Rivalry). Some notable results stand out: several countries, including Germany, Greece, Italy, Mexico, and Portugal, exhibit lower TraCE scores across all SSPs. This indicates that their observed data features are less similar to their corresponding SSP projections, when compared to most other countries in the study. Additionally, some countries deviate from the majority SSP ranking pattern. For example, Greece aligns most closely with SSP4, followed by SSP2 and SSP3, with SSP1 and SSP5 ranking the lowest. Italy aligns most with SSP3, showing strong divergence from the remaining SSPs, which have similar TraCE scores. Poland closely aligns with SSP4, followed by SSP3, with TraCE scores diverging significantly from the other SSPs.
@@
Figure 4 shows the cumulative TraCE score time series (2015-2022) for Norway, which was identified as a representative country. The TraCE score trajectories are consistently positive across all SSPs, in agreement with the expectation that they were developed as realistic scenarios in alignment with the factual historical data. Of note is the visible flattening around the year 2020, which coincides with the onset of the COVID-19 pandemic. This flattening likely occurs because the SSP projections did not anticipate the pandemic, so the observed data features deviate from these trajectories, resulting in low or negative instantaneous TraCE scores. Overall, Figure 4 indicates that SSP5 consistently ranks the highest from 2016 onwards, while other SSPs score more closely together. However, starting in mid-2021, the SSP4 and SSP1 TraCE scores begin to diverge above those of SSP2 and SSP3. With refinement, future work could correlate temporal TraCE scores with societal events and political decisions, such as legislation. Additional plots presenting the findings for Poland, as a contrasting example, are available in Appendix A.3, including a heatmap of feature importance to provide preliminary explainability of TraCE.
It must be emphasised that this study serves as a proof of concept, and requires input from experts across multiple domains to ensure safe and trustworthy implementation. This includes the selection of data features for monitoring, and their weighting, which has been equally distributed in this demonstration. Different weighting schemes will yield distinct results and should be developed in accordance with the priorities and specific questions of the user. Additionally, the data used and results obtained are contingent on the model source for SSP projections.
The utility of TraCE scores in this application lies in the capability to reconcile complex and occasionally conflicting variables into a single value. This allows experts and non-experts to quickly assess alignment with the established SSP scenarios, via an explainable method based on direction and distance in the data feature space. Visually assessing such alignment from the raw data itself can be challenging, particularly as the number of included features increases. The TraCE method is therefore useful for communication and understanding between stakeholder groups, and with refinement could aid monitoring of region sustainability against established development pathways.
@@ 5 Conclusion
TraCE provides a model-agnostic modular framework from which to assess progress over time towards an assigned goal. As demonstrated, the modularity of TraCE enables application-specific adaptation. Counterfactual target points can be defined as most appropriate, such as: model-generated counterfactuals, corpus of examples, expert-selected landmarks, or industry benchmarks.
The presented case studies involve at most 17 features. TraCE’s utility is expected to become even more evident with higher complexity scenarios which likely involve larger neural networks. In this paper we present TraCE scores in several forms: instantaneous (ICU study, Section 4.1); average and cumulative (SSP study, Section 4.2). More sophisticated methods to harness the temporal dimension could be considered after calculating TraCE scores such as quantifying instability, gradients through successive time steps, or time-dependent score weighting. The implementation of TraCE for the presented applications are for illustrative purposes, deployment and interpretation of TraCE should be guided by domain experts. Further work is required for robust implementation, including feature selection and tuning of $\lambda$.
By distilling high dimensional dynamic sequential tasks into a single value, TraCE scores enable experts and laypeople alike to quantify and better understand progress in sequential tasks.
6 Acknowledgements
We thank Thea Barnes for SQL scripts for MIMIC IV data extraction. JNC, MWLW and RSR are funded by the UKRI Turing AI Fellowship [grant number EP/V024817/1]. EAS is funded by the ARC Centre of Excellence for Automated Decision-Making and Society (project number CE200100005), funded by the Australian Government through the Australian Research Council. EFM is funded by a Google PhD Fellowship. Part of this work was done within the University of Bristol’s Machine Learning and Computer Vision (MaVi) Summer Research Program 2023.
@@ A.2 Intensive care unit outcomes
(a) Successfully discharged patient trajectory
(b) In-hospital mortality patient trajectory
Figure 6: Contrasting example patient journeys. For each, top: Instantaneous TraCE scores, higher indicates more alignment with a counterfactual. ‘Desirable’ refers to alignment with successful discharge counterfactuals, ‘Undesirable’ refers to mortality counterfactuals. TraCE is computed on the current and preceding time point, hence time point 0 is not presented. Bottom: Classifier probabilities via prediction model. Values in the legends are averages across the whole trajectory. NRFD = Not ready for discharge, RFD = ready for discharge (desirable outcome), mortality (undesirable outcome).
Figure 5(a) shows TraCE applied to another ICU patient who was successfully discharged to home. For the first two-thirds of the stay, the patient’s predicted probability of mortality was higher than for successful discharge (RFD), which is reflected by the stronger alignment with the undesirable counterfactuals (mortality) in this portion of the stay. However, the patient does recover and goes on to be successfully discharged. The TraCE score begins to increase (timepoint 7) prior to the patient’s improved health being reflected in the classifier probabilities (timepoint 8).
An unsuccessfully discharged ICU patient is shown in Figure 5(b). In this case from the TraCE score it is evident throughout the stay that the patient is deteriorating, given the consistently higher alignment with the undesirable (mortality) counterfactuals than the desirable (discharged to home) counterfactuals. This demonstrated the patient’s increasing proximity to the undesirable outcome, mortality. The increasing risk of mortality is not reflected by the classifier (Figure 5(b), bottom), which is not apparent until the patient’s final timepoint. Until this point the probability plot appeared very similar to the previously described patient (Figure 5(a)). This suggests that with refinement TraCE could provide utility, as part of a clinician’s toolkit, to support decisions and ultimately improve patient care.